Nancy Strand’s regular hourly wage rate is $32, and she receives an hourly rate of $48 for work in excess of 40 hours. During a January pay period, Nancy works 46 hours. Nancy’s federal income tax withholding is $87, and she has no voluntary deductions. Assume that the FICA tax rate is 7.65%. Prepare the employer’s journal entry to record payroll taxes for the period. Ignore unemployment taxes. (Round answers to 2 decimal places, e.g. 15.15. Credit account titles are automatically indented when amount is entered. Do not indent manually.).

What Do I enter on wiley plus? I see the solution but I'm not sure how to enter on wiley. There is only 6 boxes to enter. Please whoever helps me answer this, please answer in the format I asked. Thank You. How is this incomplete information? That is the question and format they are asking.

Date. Jan 15

account titles and explanation Debit Credit
Answer #1
Total wages = (40 * 32) + (6 * 48) = 1568
FICA tax = total wages * FICA tax rate = 1568 * 7.65% = 120
Journal Entry:
Date Account Titles and explanation Debit $ Credit $
15-Jan Employer's tax expenses 120
Payroll taxes payable 120
(being employers wages taxes made payable)